OpenStudy (anonymous):

antiderivative of 1/(x+a)(x+b)

OpenStudy (anonymous):

partial fractions... \(\large \int \frac{1}{(x+a)(x+b)}dx \) \(\large \frac{1}{(x+a)(x+b)} = \frac{A}{x+a}+\frac{B}{x+b} \) find A and B...

OpenStudy (anonymous):

Ah ok thank you!

OpenStudy (anonymous):

yw... :)